# What Factors Affect LLMs and RLLMs in Financial Question Answering?

### v1

@@ 4.3 Analysis of Agentic Frameworks
The performance gains of RLLMs within agentic frameworks primarily stem from effective agent collaboration. In RLLMs, agentic frameworks exhibit an improvement compared with prompting methods, particularly in Self-Refine and S3 Agent. This indicates that frameworks successful for LLMs can also be adapted for RLLMs. However, since RLLMs are developed through Long CoT, the performance enhancements associated with agentic frameworks, when employing longer thinking processes, are not pronounced. The benefits of these frameworks are more evident in their inherent design features, such as the reflective capabilities provided by Self-Refine and the multi-faceted thinking fostered by S3 Agent.
@@ 4.4 Analysis of Multilingual Alignment Methods
Extending the reasoning length significantly enhances the performance of LLMs. For most LLMs, employing the CLP method yields better performance than the Translate-en approach. This indicates that CLP not only aligns multiple languages to the target language but also further improves performance by strengthening the Long CoT of LLMs. Additionally, it is noteworthy that in the CLP method, the performance gain for Chinese is more pronounced than that for French. Since the initial stage alignment performance of CLP is also influenced by the inherent multilingual capabilities of the LLMs, insufficient multilingual ability may limit the performance gains achievable through CLP.
RLLMs demonstrate self-alignment capabilities for multilingual questions. In RLLMs, the effectiveness of various alignment methods, such as Translate-en and CLP, is limited. It is posited that RLLMs can achieve results comparable to those of CLP through Long CoT. This suggests that the introduction of CLP may induce overthinking in RLLMs, resulting in a decline in performance. While translation does yield some performance improvement, the enhancement is not significant.
@@ 4.5 The benefits of Long CoT cannot be offset by prompting methods and agentic frameworks
We compare the performance of Qwen-2.5-32B and DeepSeek-R1-Distill-Qwen-32B, where DeepSeek-R1-Distill-Qwen-32B is a Qwen model distilled from DeepSeek-R1, as illustrated in Figure 1. It is evident that DeepSeek-R1-Distill-Qwen-32B outperforms the base model across various methods, demonstrating an average improvement of 3.95%. This indicates that the gains from Long CoT significantly enhance the reasoning capabilities of the base model after RLLMs distillation. Consequently, it is challenging for the base model to surpass RLLMs equipped with Long CoT capabilities through advanced frameworks. While it is possible to narrow this gap using a more complex framework, we argue that this approach may lead to token consumption of LLMs approaching that of RLLMs with the same base model.
Figure 2: Statistics of average output token consumption for questions of different difficulty in LLMs/RLLMs, only counting the correct output. The line chart shows the performance of LLMs/RLLMs in the Direct method.
Model Direct Zero-shot CoT Plan-and-Solve Llama-3.1-8B-Instruct 181.50 378.34 478.02 gpt-4o-mini 198.87 395.58 493.43 gemini-1.5-flash 205.36 221.17 234.45 Qwen-2.5-32B 283.00 369.96 364.24 DeepSeek-V3 364.13 395.04 420.60 DeepSeek-R1-Distill-Qwen-32B 1,721.73 1750.14 1858.72 Qwen-3-14B 1,696.88 1701.82 1665.87
Table 3: Statistics of average correct output token consumption in current LLMs/RLLMs.
Figure 3: Performance of different scale and thinking mode in Qwen-3. The line chart in (a) shows the performance of Qwen-3 in the Direct method. The histogram in (a) shows the average correct output tokens at different scales. The histogram in (b) shows the performance of Qwen-3 in different methods.
@@ 4.6 Longer reasoning processes generally lead to better performance
To explore the relationship between output length and performance in LLMs and RLLMs, we conduct a statistical analysis of output tokens. The statistics presented in Table 3 indicate that within the same model, more robust methods require a greater number of tokens. This suggests a positive correlation between performance and token usage in the current LLMs, effective prompting techniques enhance performance by simulating a Long CoT. Statistic of Direct in Figure 2 reveals that, in LLMs, longer reasoning processes with improved performance. This suggests that sufficient cognitive processing benefits LLMs and their current limitations primarily stem from a constrained thinking process.
@@
In contrast, for RLLMs, the output of Long CoT indicates that the length of the CoT is not the primary determinant of performance. For instance, DeepSeek-R1-Distill-Qwen-32B generates an excessive number of tokens for easy questions, yet this does not lead to significant performance enhancements compared with Qwen-3-14B. In easy questions, DeepSeek-R1-Distill-Qwen-32B achieves 44.23%, Qwen-3-14B achieves 45.38% in Acc. This phenomenon may arise from the overthinking in RLLMs, which fails to yield performance gains while substantially increasing computational costs. Moving forward, dynamically adjusting CoT length according to the complexity of the input problem will represent a critical area of research.
4.7 Performance follows LLM parameter and output increase
@@
We further analyze the performance and token consumption of Qwen-3 across various scales and the thinking mode, with the final results illustrated in Figure 3. The data reveal that the performance of LLMs adheres to the scaling law Kaplan et al. (2020), larger models exhibit enhanced performance and longer output lengths. Additionally, activating the thinking mode provides a significant boost, yielding an average improvement of 35% for the Qwen-3-14B model.
Figure 4: Results of CLP on questions of different difficulty. The transparent color indicates the performance of the Direct method.
4.8 CLP significantly enhances the performance of LLMs and RLLMs when addressing complex problems
To investigate how the CLP method improves the multilingual capabilities of various LLMs and RLLMs, we analyze model performance across different difficulty levels, as illustrated in Figure 4. The results demonstrate that CLP effectively boosts model performance on challenging problems and substantially enhances multilingual proficiency. This alignment method proves to be beneficial even for RLLMs.
We posit that, for LLMs, CLP optimally enhances performance through the reasoning length and the alignment of diverse languages. For smaller LLMs, gains are observed across varying levels of difficulty. Conversely, for larger LLMs, CLP primarily improves performance on difficult tasks. While CLP can yield benefits for RLLMs through multilingual alignment, it may also lead to overthinking simpler problems, resulting in less noticeable overall gains.
@@ 4.8 CLP significantly enhances the performance of LLMs and RLLMs when addressing complex problems
Local Language English Chinese French Target Language Chinese French English French English Chinese Meta-Llama-3.1-8B-Instruct 13.78 8.05 15.57 9.97 5.12 5.76 GPT-4o-mini 29.24 27.93 27.98 25.30 12.17 12.17 Gemini-1.5-flash 30.40 30.69 27.73 28.46 9.61 12.17 Qwen-2.5-32B 32.00 31.78 31.14 31.63 14.10 14.74 DeepSeek-V3 42.38 41.79 40.87 40.38 25.64 23.71 DeepSeek-R1-Distill-Qwen-32B 35.41 36.57 35.52 35.27 19.23 21.15 Qwen-3-14B 38.60 38.24 35.76 38.92 23.07 20.51
@@
Table 4: The results of CLP. Light green presents the performance of CLP is better than the Direct method, Light red presents the performance of CLP is worse than the Direct method.
4.9 CLP must select the most effective alignment language to maximize performance
@@
We explore the performance differences across various target languages within the CLP method, with the final results presented in Table 4. The findings indicate that different target languages yield varying performance gains, and directly aligning local languages to English through CLP is often not the most effective strategy. Furthermore, different models utilize distinct optimal target languages identified by the CLP method, which may be attributed to varying language capabilities among the models. Although some researchers propose CLSP Qin et al. (2023) and AutoCAP Zhang et al. (2024a). CLSP investigates the integration of CLP and self-consistency methods to bridge the gap between different languages. AutoCAP automatically selects languages and automatic weight allocation to obtain the best language combination. This approach incurs substantial operational costs. Therefore, the selection of the best target language based on the multilingual capabilities of each model requires further exploration.
@@ 6 Conclusion
This paper is the first to explore the performance of various methods for LLMs and RLLMs in the financial question-answering task, focusing on why these methods are effective for improvement in LLMs. We conduct experiments on five LLMs and three RLLMs, revealing the critical role of Long CoT in the performance of language models. Our results indicate that the goal of these methods is to simulate Long CoT. We hope these findings deepen our understanding of the Long CoT and serve as a valuable reference in financial question-answering.
Limitations
Due to computational cost constraints, we primarily selected representative methods for our experiments. These methods may not fully reflect the performance of the majority of techniques in financial question answering. To prevent data leakage, we utilized the latest benchmark, FAMMA, for our experiments. However, we cannot guarantee that similar data exists for some questions in the internet data. The challenge of LLMs fitting data requires further exploration in future work.

### v3

@@ 4.3 Analysis of Agentic Frameworks
The performance gains of RLLMs within agentic frameworks primarily stem from effective agent collaboration. In RLLMs, agentic frameworks exhibit an improvement compared with prompting methods, particularly in S3 Agent for DeepSeek-R1-Distill-Qwen-32B. This indicates that frameworks successful for LLMs can also be adapted for RLLMs. However, since RLLMs are developed through Long CoT, the performance enhancements associated with agentic frameworks, when employing longer thinking processes, are not pronounced. The benefits of these frameworks are more evident in their inherent design features, such as the reflective capabilities provided by Self-Refine and the multi-faceted thinking fostered by S3 Agent.
@@ 4.4 Analysis of Multilingual Alignment Methods
Extending the reasoning length significantly enhances the performance of LLMs. For most LLMs, employing the CLP method yields better performance than the Translate-en approach. Additionally, the agentic framework (Self-Refine and S3 Agent) enhances the performance of LLMs to some degree. This indicates that CLP not only aligns multiple languages to the target language but also further improves performance by strengthening the Long CoT of LLMs. Since the initial stage alignment performance of CLP is also influenced by the inherent multilingual capabilities of the LLMs, insufficient multilingual ability may limit the performance gains achievable through CLP.
RLLMs demonstrate self-alignment capabilities for multilingual questions. In RLLMs, the effectiveness of various alignment methods, such as Translate-en and CLP, is limited. It is posited that RLLMs can achieve results comparable to those of CLP through Long CoT. This suggests that the introduction of CLP may induce overthinking in RLLMs, resulting in a decline in performance. While translation does yield some performance decline, the reduction is not significant.
@@ 4.5 The benefits of Long CoT cannot be offset by prompting methods and agentic frameworks
We compare the performance of Qwen-2.5-32B and DeepSeek-R1-Distill-Qwen-32B, where DeepSeek-R1-Distill-Qwen-32B is a Qwen-2.5 model distilled from DeepSeek-R1, as illustrated in Figure 1. It is evident that DeepSeek-R1-Distill-Qwen-32B outperforms the base model across various methods, demonstrating an average improvement of 7.4%. This indicates that the gains from Long CoT significantly enhance the reasoning capabilities of the base model after RLLMs distillation. Consequently, it is challenging for the base model to surpass RLLMs equipped with Long CoT capabilities through advanced frameworks. While it is possible to narrow this gap using a more complex framework, we argue that this approach may lead to token consumption of LLMs approaching that of RLLMs with the same base model.
Figure 2: Statistics of average output token consumption for questions of different difficulty in LLMs/RLLMs. The line chart shows the performance of LLMs/RLLMs in the Direct method.
Model Direct Zero-shot CoT Plan-and-Solve Meta-Llama-3.1-8B-Instruct 464.36 973.27 1131.86 gpt-4o-mini 249.77 430.41 531.83 gemini-1.5-flash 286.32 296.03 313.42 Qwen-2.5-32B 368.91 470.99 476.74 DeepSeek-V3 425.65 450.58 493.32 DeepSeek-R1-Distill-Qwen-32B 1,972.60 2006.39 2196.63 Qwen-3-14B 2,087.49 2120.97 2211.74
Table 3: Statistics of average output token consumption in current LLMs/RLLMs.
@@ 4.6 Longer reasoning processes generally lead to better performance
To explore the relationship between output length and performance in LLMs and RLLMs, we conduct a statistical analysis of output tokens. The statistics presented in Table 3 indicate that within the same model, more robust methods require a greater number of tokens. This suggests a positive correlation between performance and token usage in the current LLMs, effective prompting techniques enhance performance by simulating a Long CoT. Statistic of Direct in Figure 2 reveals that, in most LLMs, longer reasoning processes with improved performance. This suggests that sufficient cognitive processing benefits LLMs and their current limitations primarily stem from a constrained thinking process.
Figure 3: Performance of different scale and thinking mode in Qwen-3. The line chart in (a) shows Direct performance of Qwen-3, while the histogram in (a) shows average output tokens across scales. The histogram in (b) shows performance in different methods.
@@
In contrast, for RLLMs, the output of Long CoT indicates that the length of the CoT is not the primary determinant of performance. For instance, DeepSeek-R1-Distill-Qwen-32B generates an excessive number of tokens for easy questions, yet this does not lead to significant performance enhancements compared with Qwen-3-14B. In easy questions, DeepSeek-R1-Distill-Qwen-32B achieves 53.41%, Qwen-3-14B achieves 56.76% in Acc. This phenomenon may arise from the overthinking in RLLMs (Chen et al., 2024), which fails to yield performance gains while substantially increasing computational costs. Moving forward, dynamically adjusting CoT length according to the complexity of the input problem will represent a critical area of research.
4.7 Performance improves as the parameters of LLMs increase
@@
We further analyze the performance and token consumption of Qwen-3 across various scales and the thinking mode, with the final results illustrated in Figure 3. The data reveal that the performance of LLMs adheres to the scaling law (Kaplan et al., 2020), larger models exhibit enhanced performance and longer output lengths. This suggests that more powerful LLMs engage in more extensive reasoning processes, resulting in superior thinking chain capabilities.
Additionally, we examine the performance of the Qwen-3-14B model with thinking mode activated and deactivated. Activating the thinking mode yields a significant improvement, with an average increase of 16.9%. This indicates that RLLMs can achieve superior performance through extended reasoning processes.
@@
Source Language English Chinese French Target Language Chinese French English French English Chinese Meta-Llama-3.1-8B-Instruct 19.88 14.00 23.35 13.38 17.30 18.58 GPT-4o-mini 39.47 38.53 38.19 35.03 32.05 28.84 - Gemini-1.5-flash 41.14 41.50 37.22 38.92 29.48 30.76 Qwen-2.5-32B 45.57 44.99 43.30 42.82 35.25 33.97 DeepSeek-V3 62.55 60.88 61.31 59.36 57.69 53.84 DeepSeek-R1-Distill-Qwen-32B 50.79 51.45 50.85 48.90 52.56 51.28 Qwen-3-14B 57.91 57.03 53.77 58.39 51.28 50.00
@@
Table 4: The results of CLP. presents the performance of CLP is better than the Direct method, presents the performance of CLP is worse than the Direct method.
4.8 It’s important to select the optimal alignment language in CLP
We explore the performance differences across various target languages within the CLP method, with the final results presented in Table 4. The findings indicate that different target languages yield varying performance gains, and directly aligning local languages to English through CLP is often not the most effective strategy. Utilizing various language alignments has demonstrated differing impacts on performance (Wang et al., 2025b), likely attributable to LLMs’ comprehension of distinct language families.
@@ 4.8 It’s important to select the optimal alignment language in CLP
Furthermore, different models utilize distinct optimal target languages identified by the CLP method, which may be attributed to varying language capabilities among the models. Although some researchers propose CLSP (Qin et al., 2023) and AutoCAP (Zhang et al., 2024a). CLSP investigates the integration of CLP and self-consistency methods to bridge the gap between different languages. AutoCAP automatically selects languages and automatic weight allocation to obtain the best language combination. This approach incurs substantial operational costs. Therefore, the selection of the best target language based on the multilingual capabilities of each model requires further exploration.
Figure 4: Error statistics for Qwen-3-14B under different thinking modes, considering cases where thinking mode is correct but non-thinking mode fails.
4.9 Understanding the Advantages of Long CoT
To elucidate the advantages conferred by Long CoT, we utilize GPT-4o-mini to systematically analyze the error types observed in Qwen-3-14B under both thinking and non-thinking modes, as illustrated in Figure 4.
The analysis reveals that, in the absence of the thinking mode, the reasoning error constitutes the predominant source of performance degradation. In contrast, activation of the thinking mode substantially mitigates such errors, highlighting the efficacy of Long CoT in enhancing the reasoning capabilities of RLLMs. Furthermore, RLLMs demonstrate superior comprehension of problem statements compared to LLMs, as evidenced by the fact that approximately one-third of LLM errors are attributable to misinterpretation of questions and underlying assumptions. These findings collectively underscore the critical role of Long CoT in advancing the overall performance of financial question answering models.
5 Discussion
While considerable research has focused on enhancing the performance of LLMs across various domains, our experiments indicate that the key to improving LLM performance may lie in strengthening their long-term thinking capabilities, which aligns with the overarching objective of RLLMs. Consequently, we find that currently effective prompting methods and agentic frameworks for LLMs do not yield significant improvements for RLLMs. It is important to note that the agentic framework employed in our study was not specifically tailored for financial question answering. Thus, agentic frameworks that integrate RLLMs with techniques such as retrieval-augmented generation (RAG) and are well-designed for financial question answering may present a promising solution.
Furthermore, the Long CoT approach constrains the potential for enhancing performance of RLLMs through agentic frameworks. The advent of agentic reinforcement learning (Agentic RL) enables LLMs and RLLMs to function not only as standalone foundation models but also to be incorporated into complex reasoning and decision-making processes (Luo et al., 2025b; Zhang et al., 2025). Developing a suitable agentic workflow and further training LLMs and RLLMs through agentic reinforcement learning represents an intriguing avenue for improving model performance.
@@ 7 Conclusion
In this work, we systematically investigate the factors influencing the performance of LLMs and RLLMs in financial question answering. Through extensive experiments on multiple models and methods, we find that the effectiveness of prompting strategies, agentic frameworks, and multilingual alignment approaches for LLMs largely stems from their ability to simulate longer reasoning chains (Long CoT). RLLMs, equipped with inherent Long CoT capabilities, show limited improvement from these conventional methods, highlighting a performance bottleneck. Our analysis suggests that future advances for RLLMs may require more sophisticated agentic mechanisms and dynamic reasoning processes. We hope these findings deepen our understanding of the Long CoT and serve as a valuable reference in financial question-answering.
The Use of Large Language Models (LLMs)
We declare that only LLMs were utilized to polish the English of this paper.